### v1

# Spatial Association Between Near-Misses and Accident Blackspots in Sydney, Australia: A Getis-Ord $G_{i}^{*}$ Analysis

Abstract
Road safety management teams utilize on historical accident logs to identify blackspots, which are inherently rare and sparse in space and time. Near-miss events captured through vehicle telematics and transmitted in real-time by connected vehicles reveal a unique potential of prevention due to their high frequency nature and driving engagement on the road. There is currently a lack of understanding of the high potential of near-miss data in real-time to proactively detect potential risky driving areas, in advance of a fatal collision. This paper aims to spatially identify clusters of reported accidents (A) versus high-severity near-misses (High-G) within an urban environment (Sydney, Australia) and showcase how the presence of near-misses can significantly lead to future crashes in identified risky hotspots. First, by utilizing a 400m grid framework, we identify significant crash hotspots using the Getis-Ord $G_{i}^{*}$ statistical approach. Second, we employ a Bivariate Local Moran’s I (LISA) approach to assess and map the spatial concordance and discordance between official crash counts (A) and High-G counts from nearmiss data (High-G). Third, we classify areas based on their joint spatial patterns into: a) High-High (HH) as the most riskiest areas in both historical logs and nearmiss events, High-Low (HL) for high crash logs but low nearmiss records, c) Low-High (LH) for low past crash records but high nearmiss events, and d) Low-Low (LL) for safe areas. Finally, we run a feature importance ranking on all area patterns by using a contextual Point of Interest (POI) count features and we showcase which factors are the most critical to the occurrence of crash blackspots.
Index Terms:
connected vehicles, nearmisses, blackspots, statistical inference
I Introduction
I-A Background and Motivation
Road traffic accidents impose significant societal costs globally, demanding effective safety management strategies. Historically, identifying high-risk locations has predominantly relied on analysing police-reported accident data (A). While valuable, this approach is reactive, requiring accidents to occur before interventions are typically considered. Furthermore, accident data, particularly for severe incidents, can be sparse in space and time, making statistical identification of hazardous locations challenging [1].
The emergence of vehicle telematics and advanced driver-assistance systems (ADAS) has enabled the collection of vast amounts of data on driver behaviour and vehicle kinematics, including near-miss events [2]. Near-misses, defined as unsafe events where a collision is narrowly avoided, occur much more frequently than actual accidents and are considered leading indicators of underlying safety risks [3, 4]. Analysing near-miss patterns offers a proactive approach to identifying potentially hazardous locations before serious accidents occur.
Within the spectrum of near-miss data, events characterised by high severity metrics – such as large longitudinal or lateral accelerations – are hypothesised to represent situations with a higher potential for resulting in injury should a collision occur [5]. We further define this as G-Force, a measure of the acceleration during a nearmiss event (and later on as NM+G a nearmiss with a high G-force).We believe that by utilizing NM+G severity data, we can proactively refine the safety analysis by focusing on locations where the “potential consequences” of frequent conflicts are much higher.
I-B Research Gap, Objectives, and Contribution
Gap: Despite the increasing availability of near-miss frequency data, the spatial relationship between hotspots derived from near-miss severity indicators (such as high G-force events, hereafter High-G) and hotspots identified using traditional accident data (A) remains relatively unexplored, particularly at fine spatial resolutions within urban areas. To the best of our knowledge, it has never been proved whether areas exhibiting frequent high-severity near-miss events (NM+G) directly correspond spatially to areas where accidents are historically concentrated, or if NM+G events will significantly increase the occurrence of actual crashes. Establishing this link is critical for validating the use of High-G severity as a reliable spatial proxy for accident risk, enabling a more effective proactive safety management, and revealing different dimensions of underlying risk. Significant spatial discordance, for instance, might indicate areas with a high latent risk undetected by sparse accident data, or conversely, areas where underlying factors successfully mitigate the consequences of frequent severe near-misses.
Objective: Therefore, this research aims to evaluate the impact of high-severity near-miss events (High-G) on blackspot areas (A) already mapped by past historical traffic incident logs. For This study we use data provided by Compass IoT, a leading Australian startup collecting real-time data from Connected Vehicles across Australia, with a dedicated focus on Sydney, the largest city with the highest traffic incident levels. We use a 400m grid framework established for the study period (2022), and we quantitatively characterise and correlate these lagging and leading safety indicators. The specific objectives are to:
  1. 1.
    Identify statistically significant spatial clusters (hotspots/coldspots) of reported accidents (A) via the Getis-Ord $G_{i}^{*}$ statistic on the 400m grid.
  2. 2.
    Aggregate High-G event data onto the same 400m grid framework for fine resolution mapping.
  3. 3.
    Quantify and map the local spatial correlation between aggregated crash and High-G counts using Bivariate Moran’s I (LISA) to identify distinct concordance/discordance patterns (HH, LL, HL, LH).
  4. 4.
    Characterise the identified LISA pattern areas using Point of Interest (POI) counts and assess the contribution of these environmental features in differentiating key spatial risk profiles (e.g., HH vs. LL).
This evaluation of local spatial correlation patterns (via LISA) between leading (High-G) and lagging (A) indicators provides a nuanced understanding of potential versus realized traffic risk across diverse urban settings. Characterizing these patterns with POI data further elucidates the limited role of such static environmental context in explaining risk variations. The findings directly inform the practical application and limitations of using telematics-derived severity data for network screening and targeted safety interventions.
II Related Works
Road safety analysis has traditionally centered on reactive approaches, primarily identifying high-risk locations, often termed hotspots or blackspots, based on historical police-reported accident data [1, 6]. Methodologies evolved from simple frequency rankings or accident rate calculations to more sophisticated spatial statistical techniques deployed within Geographic Information Systems (GIS) [7]. Prominent methods include Kernel Density Estimation (KDE) for visualizing density [8] and spatial autocorrelation analyses, such as Moran’s I for assessing global clustering [9] and Local Indicators of Spatial Association (LISA) like Getis-Ord $G_{i}^{*}$ and Local Moran’s I for pinpointing statistically significant local clusters of high (hotspots) or low (coldspots) incident counts [9, 10, 11]. However, the fundamental limitations of this approach are well-documented: its reactive nature (requiring crashes to occur first), the relative rarity and potential underreporting or inaccuracy of official crash data, and the ethical concerns associated with waiting for harm [1, 12, 13].
These limitations spurred a significant shift towards proactive methodologies leveraging Surrogate Safety Measures (SSMs) [12]. The foundation lies in the Traffic Conflict Technique (TCT), formalized decades ago to systematically observe near-miss events [3, 14]. A near-miss or traffic conflict is generally defined as an interaction necessitating an evasive maneuver to avoid a collision [12, 3]. These non-crash events occur far more frequently than actual crashes, providing statistically richer datasets for analyzing underlying traffic risks and evaluating countermeasures without relying on sparse crash data [12, 4]. Consequently, research has focused on developing and applying various SSM indicators derived from detailed observational or sensor data. Common indicators include temporal measures like Time-to-Collision (TTC) and Post-Encroachment Time (PET), deceleration requirements like DRAC, and kinematic indicators such as speed, lateral deviation, or harsh events (e.g., rapid braking/acceleration) [12, 4]. Data for SSM calculation is increasingly sourced from video analytics platforms employing computer vision and AI [15, 16, 17], instrumented vehicles in Naturalistic Driving Studies (NDS) [2, 18], smartphone sensors [4], and Connected Vehicle (CV) data streams (e.g., Basic Safety Messages) [4, 19].
Analyzing crash causation to develop effective safety countermeasures is essential, yet hindered by the low frequency of actual crash events, particularly within rich datasets from Naturalistic Driving Studies (NDS) [20]. NDS provides detailed real-world driving data but typically captures few crashes relative to the volume of driving. This data scarcity necessitates the use of Surrogate Safety Measures (SSMs) – observable events, like traffic conflicts or near-crashes, thought to be correlated with crash risk [4]. However, establishing the validity of SSMs as reliable proxies for crash risk remains a significant and ongoing challenge in traffic safety research [21].
Near-crashes, commonly defined as events requiring a rapid evasive maneuver to avoid a collision, are frequently used SSMs, especially in NDS analysis. Their application often relies on the “causal continuum” hypothesis, which posits that near-crashes and crashes arise from largely similar or identical causal factors. This assumption is considered plausible and supported by reported correlations between conflict/near-crash frequency and historical crash rates [4]. Foundational research by Guo et al. [22] provided critical support for this approach by demonstrating that near-crashes often exhibit kinematic signatures similar to crashes and appear to share common underlying causal mechanisms or contributing factors. These findings established near-crashes as viable and effective surrogates, allowing researchers to analyze the more abundant near-crash data to understand risk and causation.
Extreme Value Theory (EVT) offers a promising statistical framework to formally link the distribution of frequent surrogate events to the probability of rare, extreme crash events [12, 5, 23], though its application requires careful consideration of assumptions and data quality [5]. Furthermore, the effectiveness and interpretation of SSMs are highly context-dependent, influenced by factors like traffic composition (homogeneous vs. heterogeneous), road geometry (intersections, curves), and environmental conditions, necessitating context-specific indicator selection and threshold calibration [12, 13, 24]. Spatial analysis of near-miss frequency hotspots has become an area of growing interest for identifying general conflict-prone areas [17].
While near-miss frequency indicates the prevalence of conflicts, metrics reflecting near-miss severity, such as high G-force events (often denoted NM+G) captured by inertial sensors in vehicles or smartphones, offer outlook into the potential consequence or danger level of these interactions [4]. High G-force readings signify rapid changes in velocity indicative of harsh braking or abrupt evasive manoeuvres, which may be associated with more dangerous situations compared to less severe conflicts. However, research directly investigating the relationship between the severity of surrogate events and the severity of actual crashes is less developed than frequency-based comparisons. Specifically, there is limited work that directly compares the spatial patterns of hotspots derived from near-miss severity indicators (like NM+G) with hotspots identified using crash severity data. This study aims to address this specific gap by employing established spatial statistical methods (Getis-Ord $G_{i}^{*}$) to explicitly examine the concordance and discordance between hotspots identified using high G-force near-miss data and those identified using historical crash data, stratified by crash severity. This comparison seeks to clarify the utility of severity-based near-miss data as a spatial proxy for realized accident risk, particularly concerning more severe crash outcomes.
III Case study
This study integrates three key spatial datasets for the Sydney Greater Metropolitan Area, focusing on the year 2022:
  • Road Accidents: Locations of reported traffic crashes extracted for the year 2022 from a dataset (originally covering Jan 2017 - Jul 2022) sourced from Transport for New South Wales (TfNSW) (n = 3,658 points).
  • High G-Force Near Misses: Locations identified from vehicle trajectory data provided by Compass IoT for 2022. These represent the point of maximum G-force recorded within vehicle trajectories where a near-miss event was detected (n = 24,137 points).
  • Points of Interest (POIs): Geographic locations of various amenities, shops, and other features extracted from OpenStreetMap (OSM) data to provide an environmental context.
IV Statistical Analysis Methodology
The core of this study involved a quantitative spatial analysis of reported crashes (A, $n=$3638$$) and high G-force events (High-G, $n=$23999$$) within Sydney (mapped for year 2022), using data aggregated onto a uniform 400m grid. The methodology focused on identifying statistically significant crash hotspots, analyzing the local spatial correlation between crashes and High-G events, and characterizing the resulting spatial patterns using Point of Interest (POI) data. Key stages that have been applied are the following:
IV-A Spatial Framework and Data Aggregation
To facilitate our area-based analysis, a uniform 400m x 400m grid was established across the study region (Projected CRS: EPSG:32756), resulting in $38824$ cells. The size of the grid has been selected after conducting a sensitivity analysis of the best grid size to provide granular view into risky manoeuvres occurring even across smaller secondary roads/nieghbourhood alleys. Data of the reported crashes and High-G events were spatially aggregated onto this grid, yielding cell-level counts for each variable (crash_count, highg_count). These aggregated counts formed the primary input for the spatial statistical analyses described below. Point-based POI data was processed separately for subsequent feature generation (see IV-C).
IV-B Crash Hotspot Identification and Bivariate Correlation
Spatial clustering and correlation were assessed using established geospatial statistics applied to the aggregated grid data:
  • Defining Spatial Relationships: A Spatial Weights Matrix (SWM) based on Queen contiguity was constructed to formally define the neighborhood structure and spatial influence between adjacent grid cells.
  • Crash Hotspot Detection (Gi*): The Getis-Ord $G_{i}^{*}$ statistic, a local indicator of spatial association, was calculated for each grid cell based solely on its crash_count relative to its neighbors (defined by the SWM). Significance testing ($p<0.10$, $p<0.05$, $p<0.01$) using permutation inference identified statistically significant crash hot spots (high-crash clusters) and cold spots (low-crash clusters).
  • Bivariate Spatial Correlation (LISA): Bivariate Local Moran’s I (LISA) analysis was employed to quantify and map the local spatial correlation between crash_count and highg_count. This identified cells exhibiting statistically significant ($p<0.05$) spatial patterns: High Crash-High HighG (HH), Low Crash-Low HighG (LL), High Crash-Low HighG (HL), and Low Crash-High HighG (LH), revealing areas of concordance and discordance between the two indicators.
IV-C Characterization of LISA Patterns using POI Features
Following the identification of distinct spatial correlation patterns via LISA, a characterization focused on the environmental context provided by Points of Interest: POI data (sydney_pois_filtered.csv), containing point locations and types, was spatially joined with the analysis grid. For each grid cell, the number of POIs of each distinct type falling within its boundary was counted using geopandas, generating POI count features (e.g., poi_type_Park, poi_type_School).
IV-D Spatial Clustering of Crash Events
The Getis-Ord $G_{i}^{*}$ statistic was employed to identify statistically significant spatial clusters of reported crashes (A) across the 400m grid cells covering the Sydney study area (circa 2022). This identifies localized concentrations significantly higher (hotspots) or lower (coldspots) than expected given the overall spatial distribution. The resulting spatial pattern of crash hotspots and coldspots is illustrated in Fig. 1.
Figure 1: Spatial distribution of statistically significant road accidents clusters on the 400m grid, based on Getis-Ord $G_{i}^{*}$ analysis for 2022. Red/salmon areas indicate hotspots (significantly high concentration), blue areas indicate coldspots (significantly low concentration), and grey areas represent locations with no statistically significant clustering at the $p<0.05$ level.
Local Spatial Correlation between Crash Counts and High-G Event Counts: To investigate the local relationship between the frequency of reported crashes and the frequency of high G-force (High-G) events, a Bivariate Local Moran’s I (LISA) analysis was performed on the 400m grid cell data. This analysis assesses whether the spatial pattern of crash counts in a given cell is significantly correlated with the spatial pattern of High-G event counts in its neighborhood (using Queen contiguity). The geographical distribution of these local spatial correlations is depicted in 2.
Figure 2: Map of Bivariate Local Moran’s I (LISA) results on the 400m grid, showing the spatial correlation between crash counts and High-G event counts per cell. Colors represent the type of statistically significant ($p<0.05$) local correlation: High-High (Red: high crashes, high High-G), Low-Low (Blue: low crashes, low High-G), High-Low (Pink: high crashes, low High-G), Low-High (Light Blue: low crashes, high High-G). Grey areas indicate no significant local spatial correlation. (Note: Analysis log saved this plot as crash_highg_correlation.pdf)
V Results
V-A Spatial Analysis of Crash and High-G Event Clustering (400m Grid)
The spatial distribution and correlation of reported crashes ($n=$3638$$) and high G-force (High-G) events ($n=$23999$$) were analyzed using a 400m x 400m square grid resolution, encompassing $38824$ cells across the Sydney study area. Initial analysis using Global Moran’s I confirmed significant positive spatial autocorrelation (clustering) for both crash counts (Moran’s I = $0.2080$, $p=0.001$) and High-G event counts (Moran’s I = $0.2586$, $p=0.001$). This indicates an overall tendency for both phenomena to cluster spatially rather than being randomly distributed.
Furthermore, the overall spatial relationship between the two variables was assessed using Global Bivariate Moran’s I, which revealed a significant positive association (Global Bivariate Moran’s I = $0.1654$, $p=0.001$). This suggests that, on average, areas with high crash counts tend to be spatially close to areas with high High-G event counts across the study region.
To explore the specific local patterns of spatial association between crash frequency and the near-miss proxy (High-G events), Bivariate Local Moran’s I (LISA) was used (using $p<0.05$ for significance). This method classified the grid cells based on their significant local spatial relationship patterns.
V-B Identification and Classification of Near-Miss Hotspots
This analysis investigates the spatial relationship between potential traffic conflicts, represented by high G-force events (High-G, $n=23,999$), and realized harm, represented by reported crash locations (Crashes, $n=3,638$). The study area was divided into a regular grid with 400m x 400m cells.
To understand the spatial concordance or discordance between these indicators, a Bivariate Local Moran’s I (LISA) analysis was conducted. This spatially correlates the counts of High-G events with the counts of reported Crashes within the same grid cell system (using Queen contiguity weights). The analysis identifies locations where the local concentration of High-G events significantly aligns (or fails to align) with the local concentration of Crashes. Based on the LISA results (significant at $p<0.05$), grid cells were classified into distinct spatial relationship categories. These classifications reveal distinct spatial patterns based on the Bivariate Local Moran’s I (LISA) analysis comparing crash counts and high G-force (High-G) event counts within 400m grid cells (Table I):
  • HH (High Crash-High High-G): Represents 825 grid cells where high counts of potential conflicts (High-G events) spatially coincide with high counts of realized harm (Crashes). These areas strongly indicate locations where risky driving maneuvers are frequent and may directly translate into reported crashes. These are the primary co-located hotspots.
  • HL (High Crash-Low High-G): Represents 91 grid cells with historically high crash counts that are *not* statistically associated with high counts of the measured High-G events in their vicinity. This suggests that crash risk in these locations might be driven by factors other than those frequently captured by the High-G metric (e.g., specific static hazards, complex intersection designs not inducing frequent harsh braking/swerving, reporting biases, or different types of risky behavior).
  • LH (Low Crash-High High-G): Represents a substantial number of areas (2681 grid cells) characterized by statistically significant high counts of potential conflicts (High-G events) but low counts of actual reported crashes. These ’Emerging Risk’ or ’Near-Miss Hotspot’ locations are of particular interest. They suggest the presence of frequent risky maneuvers or situations, but perhaps mitigating factors (e.g., effective road design allowing recovery, lower speeds, successful evasive actions, under-reporting of minor crashes) currently prevent these from translating into high numbers of reported crashes. These areas warrant proactive monitoring.
  • LL (Low Crash-Low High-G): Represents areas where low potential conflict (High-G) and low realized harm (Crashes) would theoretically coincide, likely indicating baseline safer conditions or lower exposure areas. However, the analysis found no grid cells (0 cells) exhibiting a statistically significant Low-Low spatial relationship at the $p<0.05$ level. While many areas likely have low counts of both, they don’t form statistically significant spatial clusters of low-low association according to this specific bivariate test.
  • Not Significant: The majority of grid cells (35227) did not show a statistically significant spatial correlation ($p\geq 0.05$) between local crash counts and neighbouring High-G counts (or vice-versa) under the Bivariate LISA test.
These local patterns highlight specific areas of concordance (HH) and discordance (HL, LH) between historical crash data and the High-G near-miss indicator.
TABLE I: Classification and Size of Grid Cells Based on Bivariate Spatial Correlation (LISA, $p<0.05$) between Crash Counts and High G-Force Event Counts.
Spatial Relationship LISA Classification Number of Cells (n)
High Crash & High High-G HH (High Crash-High HighG) 825
High Crash & Low High-G HL (High Crash-Low HighG) 91
Low Crash & High High-G LH (Low Crash-High HighG) 2681
Low Crash & Low High-G LL (Low Crash-Low HighG) 0*
Not Spatially Correlated Not Significant (LISA) 35227
* No grid cells showed a statistically significant Low-Low spatial relationship at $p<0.05$.
Comparing these different spatial clusters, particularly the high number of LH cells (2681) versus HH cells (825), suggests that while High-G events are spatially associated with crashes in many hotspots (HH), there are even more areas where frequent near-misses occur without a corresponding high crash history (LH). This highlights the potential of High-G data for proactive safety analysis, identifying areas of concern before they become crash blackspots, and potentially revealing locations where safety interventions or specific road characteristics are effectively mitigating crash outcomes despite frequent risky events. Further analysis of the characteristics differentiating HH, HL, and LH areas is crucial.
Interpretation of Mann-Whitney U Test Results
Findings from the Mann-Whitney U tests conducted to compare the prevalence of various Points of Interest (POIs) between grid cells classified as ’High Crash-High HighG’ (Clusters) and those classified as ’Low Crash-Low HighG’ (Outliers) (see Table II). The tests aimed to identify statistically significant differences in the distribution of POI counts between these two area types, using a significance level ($\alpha$) of 0.05.
It is important to note that the Points of Interest (POIs) analyzed in this study originate from two primary sources: OpenStreetMap (OSM) data (specifically using tags like amenity, shop, tourism, etc) and a dedicated traffic light dataset provided by Transport for NSW.
TABLE II: Mann-Whitney U Test Results: POI Counts in High Crash-High HighG vs Low Crash-Low HighG Areas
POI Type U Statistic p-value Mean Outliers (LL) Mean Clusters (HH) Significant ($\alpha=0.05$)
appliance 1109934.0 0.001791 0.000000 0.003636 True
bed 1109934.0 0.001791 0.000000 0.006061 True
viewpoint 1122576.0 0.001928 0.014174 0.027879 True
radiotechnics 1108593.5 0.010795 0.000000 0.002424 True
flooring 1108593.5 0.010795 0.000000 0.002424 True
trophy 1108593.5 0.010795 0.000000 0.002424 True
lighting 1108593.5 0.010795 0.000000 0.002424 True
prep_school 1108593.5 0.010795 0.000000 0.002424 True
baby_goods 1108593.5 0.010795 0.000000 0.002424 True
fabric 1110449.5 0.012690 0.000746 0.004848 True
health_food 1109521.5 0.015208 0.000373 0.003636 True
pawnbroker 1109521.5 0.015208 0.000373 0.003636 True
traffic_signals 1076733.0 0.017472 0.441999 0.303030 True
waste_basket 1123232.0 0.022249 0.082432 0.136970 True
bench 1128707.0 0.030349 0.303991 0.449697 True
Key Observations:
  1. 1.
    Distinct POI Profiles: The results strongly indicate that the environmental characteristics, as represented by POI types, differ significantly between the High Crash-High HighG and Low Crash-Low HighG areas.
  2. 2.
    POIs More Prevalent in High Crash-High HighG Areas: A notable number of POI types were found to be significantly more common in the High-High cluster areas. These include:
    • Retail and Commercial Services: A variety of retail outlets (appliance, bed, radiotechnics, flooring, trophy, lighting, baby_goods, fabric, health_food, pawnbroker, carpet) showed significantly higher mean counts in High-High areas. For many of these, the mean count in Low-Low areas was zero or near-zero, suggesting their presence is a stronger characteristic of the High-High areas within this dataset.
    • Public Amenities and Features: Features such as viewpoint, waste_basket, and bench were also significantly more prevalent in High-High areas. The higher density of benches and viewpoints might suggest areas with higher pedestrian activity or specific urban design features.
    • Educational Facilities: prep_school counts were significantly higher in the High-High areas.
  3. 3.
    POIs Less Prevalent in High Crash-High HighG Areas: Perhaps the most striking finding in this category is the significantly lower prevalence of traffic_signals in High-High areas compared to Low-Low areas (p=0.017). This suggests that the High-High areas, despite having more crashes and High-G events, might be less characterized by major, signalized intersections compared to the Low-Low areas used in this comparison. This could point towards differences in road network hierarchy, traffic control strategies, or potentially higher prevalence of unsignalized intersections or different road types (e.g., mid-block segments) contributing to the High-High classification.
  4. 4.
    No Significant Difference: Several POI types, including dive_centre, music_school, safety_equipment, grave_yard, arts_centre, handwashing, and tea, did not show a statistically significant difference between the two groups. These POIs were generally rare in both area types, as indicated by their low mean counts.
VI Conclusion
This study investigated the spatial relationship between reported road crashes and a high-severity near-miss proxy (High-G events) in Sydney using Bivariate Local Moran’s I (LISA) over a 400m grid. The analysis confirmed significant spatial clustering of crashes and revealed distinct local patterns of association between the two indicators. Concordant hotspot areas were identified where high crash rates coincided with statistically high neighbouring High-G rates (HH pattern, $825$ cells). However, no areas showed a statistically significant pattern where both indicators were concurrently low (LL pattern, 00 significant cells found at $p<0.05$). Crucially, discordant areas were also found: locations with high crashes but statistically low neighbouring High-G events (HL pattern, $91$ cells), and a substantial number of locations ($2681$ cells) with low crashes despite statistically high neighbouring High-G events (LH pattern).
These findings demonstrate that while High-G events have spatial distribution does not perfectly mirror that of reported crashes. Characterizing the different pattern areas (HH, HL, LH) using Point of Interest (POI) data revealed significant differences in land use context (e.g., HH areas showing different POI profiles compared to HL or LH areas, potentially highlighting varying environmental contributors to risk). The discordant patterns are particularly valuable: HL areas ($91$) may point to crash causes not captured well by the High-G metric, while the numerous LH areas ($2681$) suggest locations with frequent risky events where crashes are currently mitigated or under-reported, warranting proactive investigation and monitoring. Limitations include data accuracy, the specific definition and proxy nature of High-G events, and potential Modifiable Areal Unit Problem (MAUP) effects from the grid analysis. Future work should incorporate dynamic traffic flow and detailed infrastructure data to better explain these complex spatial patterns and advance proactive road safety strategies.
Future research should focus on incorporating dynamic traffic variables, detailed infrastructure data, and potentially driver behavior information to build more comprehensive models explaining the observed spatial patterns of concordance and discordance. Network-based analyses and qualitative case studies of specific HL and LH locations could provide further findings into the factors mitigating or exacerbating risk. Ultimately, a multi-faceted approach combining leading and lagging indicators with rich contextual data is essential for advancing proactive road safety management.
VII ACKNOWLEDGEMENTS
We thank Compass IoT for the data and support provided for this study. This work has been funded by the UTS Jenny Edwards Fellowship awarded in 2025 to Assoc. Prof. Adriana-Simona Mihaita for conducting research in risky driver behaviour identification.